OpenStudy (goformit100):

Mr. Johnson was to earn £ 300 and a free holiday for seven weeks' work. He worked for only 4 weeks and earned £ 30 and a free holiday. What was the value of the holiday? A. £ 300 B. £ 330 C. £ 360 D. £ 420

OpenStudy (goformit100):

@Eyad

OpenStudy (anonymous):

lol give me a break ! xD

OpenStudy (anonymous):

Show your work :P

OpenStudy (goformit100):

sir tyhe questions I am posting today are simply very hard for me. I am unable to think on it :( So asking you to reply

OpenStudy (anonymous):

Let the value of the holiday =x then to pay for seven weeks work = 300$ +x so ,(300+x)/7 * 4 = 30+x Solve for x :P

OpenStudy (anonymous):

think of it as a system of equations 7weeks+1free=300 4weeks+1free=30 no subtract the second from the first to yield 3weeks=270 no solve for the amount one week is worth 1week=270/3=90 no plug 90 back into one of the equations to find what free is worth 4(90)+1free=30 360+1free=30 1free=30-260=-330 that means that instead of getting paid 330 you have to pay 330 for a free day, which means you value a free day at 330